# Focus On This, Not That! Steering LLMs With Adaptive Feature Specification

### v1

@@ Abstract
Despite the success of Instruction Tuning (IT) in training large language models (LLMs) to perform arbitrary user-specified tasks, these models often still leverage spurious or biased features learned from their training data, leading to undesired behaviours when deploying them in new contexts. In this work, we introduce Focus Instruction Tuning (FIT), which trains LLMs to condition their responses by “focusing on” specific features whilst ignoring others, leading to different behaviours based on what features are specified. Across several experimental settings, we show that focus-tuned models can be adaptively steered by focusing on different features at inference-time: for instance, robustness can be improved by focusing on task-causal features and ignoring spurious features, and social bias can be mitigated by ignoring demographic categories. Furthermore, FIT can steer behaviour in new contexts, generalising under distribution shift and to new unseen features at inference time, and thereby facilitating more robust, fair, and controllable LLM applications in real-world environments.
@@ 1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised form of supervised fine-tuning (SFT), has become an essential step in the process of developing effective instruction-following large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training to perform next token prediction allows LLMs to extract common patterns and knowledge from large text corpora, IT fine-tunes these models on input-output pairs complemented by natural-language task instructions, teaching them to perform open-ended language-based tasks given instructions (Huang et al., 2023). However, despite the improvements observed in zero-shot generalisation from IT, recent studies suggest that some of these gains may be superficial, stemming from the models’ ability to learn task template formats or spurious input/output correlations rather than a more generalisable instruction-following capability (Kung & Peng, 2023; Ghosh et al., 2024). As a result, LLMs may fail to generalise to new contexts where the same templates or spurious correlations are not present (Kung & Peng, 2023).
To address these limitations, we propose Focus Instruction Tuning (FIT), an extension of traditional IT where LLMs are also fine-tuned with respect to an instruction regarding which features to “focus on” or “not focus on”. Training LLMs to condition responses on the provided focus specification, where responses to the same input differ based on the specified features, allows end users to dynamically steer model behaviour by indicating which features should play a role or be ignored in performing a task. For instance, in Figure 1, we show how FIT can be used to steer a model to ignore gender stereotypes and focus instead on task-relevant information in order to correctly solve a question-answering task. In our experiments, we demonstrate that FIT is effective at steering models to ignore known spurious features while focusing on causal (task-relevant) features (see Section 4) across a variety of features and tasks such as sentiment analysis and natural language inference, and can be used to mitigate social bias by ignoring demographic features in question-answering. We find that FIT is also robust to distribution shifts over feature values, and can generalise to new, held-out features not encountered during training. In sum, our primary contributions are as follows:
@@
    We introduce Focus Instruction Tuning (FIT), a method that allows users to flexibly and dynamically specify what features a model should or should not focus on when performing a task at inference time. FIT enables practitioners to incorporate domain-specific knowledge about causal, spurious, or bias-relevant features in order to steer models according to the desired feature specification.
@@
    We show that FIT generalises with respect to (a) distribution shift over feature values and (b) focusing on new features not seen during training.
Figure 1: Focus Instruction Tuning (FIT). In the example above, a model that is simply Instruction Tuned would follow biases present in the training set. On the other hand, a FIT model can dynamically focus or ignore specific features of the input. Here we report a sample from BBQ (Parrish et al., 2022) (a question- answering dataset designed to elicit incorrect answers from models that perform the task by resorting to socially biased stereotypes). Both a man and a woman are present at an engineering candidate’s interview. When asked who made the technical question, due to the pre-training and instruction tuning biases, a model could respond the man did, despite the conflicting evidence present in the context. A focus-tuned model can ignore the gender feature and focus instead on the interview content.
@@ 2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset – i.e., input features that are correlated with outputs in the training distribution, but are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold Wang et al. (2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Xiao et al., 2021; Venkataramani et al., 2024; Arjovsky et al., 2019) or texture (Geirhos et al., 2018; Baker et al., 2018), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024b; Borkan et al., 2019). For instance, the token SPIELBERG is spuriously correlated with positive sentiment in datasets like SST-2 (Socher et al., 2013b), meaning that models trained on SST-2 may learn to predict sentiment by leveraging these spurious features instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-causal features undermines the robustness of models in generalising to distribution shift.
A variety of approaches have been explored to detect and mitigate the effects of spurious feature learning, particularly under distribution shifts. Traditional approaches include prompt engineering (Sun et al., 2024b), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), and directly incorporating causal inference strategies (Wang & Culotta, 2020; 2021; Udomcharoenchaikit et al., 2022). Substantial work in mechanistic interpretability has also aimed to discover models’ representation and use of task-causal or spurious features: for instance, causal probing (which trains probing classifiers to recognise and modify supervised feature representations encoded by foundation models; see Belinkov, 2022; Canby et al., 2024; Davies & Khakzar, 2024) has been used to study how models leverage causal versus spurious features features in the context of a given task (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating their use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, concept removal methods locate and manipulate supervised feature representations corresponding to bias features encoded by foundation models in order to remove these features (Ravfogel et al., 2020; 2022; 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
@@ 2.2 Controlling LLMs
Refocusing LLMs. Refocusing LLMs. Several methods have been proposed to better control instruction-tuned models both during and after training. Llama-Guard (Inan et al., 2023) fine-tunes LLMs to detect predefined risk features in inputs and outputs based on a user-specified taxonomy, such as identifying sexual content in inappropriate contexts. JsonTuning (Gao et al., 2023) enhances traditional instruction tuning by enforcing structured input and output formats in JSON, clarifying task requirements and reducing sensitivity to paraphrasing (Sun et al., 2024a). In contrast, Focus Instruction Tuning (FIT), as introduced in this work, provides a more flexible and powerful approach. While Llama-Guard operates only post-training and is limited to the safety domain, FIT enables fine-grained control both during and after training, conditioning models on a broader range of features across arbitrary domains via natural-language specifications. Moreover, unlike JsonTuning, which is restricted to enforcing output structure, FIT allows users to specify input features, enabling the model to ignore spurious correlations or highlight task-relevant attributes.
@@ 3 Methodology
Let $\mathcal{F}$ denote the set of possible features (e.g., specific keywords, sentiment, verb tense, demographic information, etc.) that the model can be instructed to focus on or ignore when generating responses. We consider a set of natural language instructions to focus or rule out specified features in $\mathcal{F}$ which we term the focus instruction set $\mathcal{I}_{\text{focus}}$. [^1] Examples of focus instructions specified in natural language include: “Make sentiment the central factor in your decision” and “Base your prediction solely on the presence of keywords. Exclude the logical relationship between the premise and hypothesis”. Explicitly, we define $\mathcal{I}_{\text{focus}}$ as
@@
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt in order to aid the model in learning the underlying task as well as the ability to refocus its attention on specified features during FIT.
@@
FIT Training. Consider a classification task [^2] For simplicity, we focus on classification here; but FIT is also applicable to generative tasks, as we show in our question-answering experiments. with finite label space $\mathcal{Y}$, where a single causal feature $C\in\mathcal{F}$ is fully predictive of label $y\in\mathcal{Y}$ given input $x$ at both training time and under distribution shift (Koh et al., 2021). We also consider spurious features $S\in\mathcal{S}\subseteq\mathcal{F}$ from a subset of spurious features $\mathcal{S}$, where feature values [^3] Note that we use uppercase to denote features, and lowercase to denote specific values of features. $s\in\text{Image}(S)$ for some spurious feature $S\in\mathcal{S}$ correlate with a label $y_{s}\in\mathcal{Y}$, where this correlation may change under distribution shift (Ming et al., 2022). Finally, we define $\mathcal{F}$ as the set of features that may be included in focus instructions during training, consisting of the causal feature and the set of spurious features =$\mathcal{F}=\{C\}\cup\mathcal{S}$.
For a sample $(x,y)\sim p_{\text{data}}(x,y)$, we specify the focus label $y_{\text{focus}}=y_{\text{focus}}(x,y,I_{\text{focus}})\in\mathcal{Y}$ that depends on the ground truth label $y$ and focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. Intuitively, we define focus label $y_{\text{focus}}$ as $y_{\text{focus}}=y$ when either no focus features are specified (i.e., using the empty focus instruction), when the focus is on the underlying causal feature $C$, or when ignoring a spurious feature $S$; but when either the focus is on a spurious feature or the causal feature is ignored, $y_{\text{focus}}$ is defined as the label spuriously correlated with a particular value of the spurious feature present in input $x$. This changing target $y_{\text{focus}}$ trains the model to learn to adjust its responses based on specified features. Formally, we define $y_{\text{focus}}$ as:
@@
(Defining $\rho_{\text{spurious}}$). Let $S\in\mathcal{S}\subseteq\mathcal{F}$ denote a spurious feature. Suppose that a value of $S$, say $s$, is spuriously correlated with label $y_{s}$. Then we define $\rho_{\text{spurious}}(s)$ as
$\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|X,S=s,s\in X)$ (4)
for some dataset sample $(x,y)\in\mathcal{D}$, where $S\in X$ denotes the presence of feature $S$ in example $X$.
By varying $\rho_{\text{spurious}}(s)$, we can control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well as causal features.
Given a task with $N$ classes, we require $\rho_{\text{spurious}}=1/N$ within the training set, ensuring that the underlying label distribution, $p(y|I,I_{\text{focus}},x)$, is of maximum entropy when focusing on spurious features. This allows the model to better distinguish between causal and spurious features, as effectively minimising Equation 3 would require the model to make predictions without relying on the underlying causal feature when its attention is specified to focus on on spurious features. A more detailed exploration of this setting of $\rho_{\text{spurious}}$ during training can be found in Section A.1.
@@
We further evaluate FIT under distribution shifts, where the specific values taken by spurious features do not overlap between the training and test sets, by introducing one additional test set:
  • $\mathcal{D}_{\text{shift}}(\rho_{\text{spurious}})$: Test datasets where the spurious feature values are distinct from those within the training set.
We evaluate over these datasets specifically on our SMNLI datset (c.f. Section 4.2).
@@ 4 Experiments
Baselines. We consider two baselines to compare against FIT: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline follows the same setup as the FIT method, but without the focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions in FIT. The few-shot baseline involves using 5 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. Further details of baselines can be found in Section A.4.

### v2

@@ Abstract
Despite the success of Instruction Tuning (IT) in training large language models (LLMs) to perform arbitrary user-specified tasks, these models often still leverage spurious or biased features learned from their training data, leading to undesired behaviours when deploying them in new contexts. In this work, we introduce Focus Instruction Tuning (FIT), which trains LLMs to condition their responses by focusing on specific features whilst ignoring others, leading to different behaviours based on what features are specified. Across several experimental settings, we show that focus-tuned models can be adaptively steered by focusing on different features at inference-time: for instance, robustness can be improved by focusing on core task features and ignoring spurious features, and social bias can be mitigated by ignoring demographic categories. Furthermore, FIT can steer behaviour in new contexts, generalising under distribution shift and to new unseen features at inference time, and thereby facilitating more robust, fair, and controllable LLM applications in real-world environments.
@@ 1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised form of supervised fine-tuning (SFT), has become an essential step in the process of developing effective instruction-following large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training to perform next token prediction allows LLMs to extract common patterns and knowledge from large text corpora, IT fine-tunes these models on input-output pairs complemented by natural-language task instructions, teaching them to perform open-ended language-based tasks given instructions (Huang et al., 2023).
However, despite the improvements observed in zero-shot generalisation from IT, recent studies suggest that some of these gains may be superficial, stemming from the models’ ability to learn task template formats or spurious input/output correlations rather than a more generalisable instruction-following capability (Kung & Peng, 2023; Ghosh et al., 2024). As a result, LLMs may fail to generalise to new contexts where the same templates or spurious correlations are not present (Kung & Peng, 2023). Furthermore, fine-tuning can inadvertently lead to safety misalignment, where models lose alignment with desired objectives and become more prone to generating harmful or undesirable outputs (Qi et al., 2023). This motivates the need for methods that produce adaptive models at inference time, enabling users to steer model behaviour based on evolving requirements or safety considerations.
To address these limitations, we propose Focus Instruction Tuning (FIT), an extension of traditional IT where LLMs are fine-tuned with respect to an instruction specifying which features to focus on” or not focus on.” By training LLMs to condition responses on these focus specifications—enabling responses to the same input to differ based on the specified features—FIT allows end users to dynamically steer model behavior. This capability not only provides precise control over which features play a role or are ignored in performing a task but also adds an important layer of explainability. For instance, in Figure 1, we demonstrate how FIT can steer a model to ignore gender stereotypes and focus on task-relevant information, enabling it to correctly solve a question-answering task.
In our experiments, we demonstrate that FIT enables precise control over model behavior by steering it to focus on task-relevant features while ignoring features specified as irrelevant or spurious (see Section 4). This flexibility allows FIT to address challenges such as mitigating the influence of demographic features in question-answering tasks, while also proving effective across diverse NLP tasks like sentiment analysis and natural language inference. Furthermore, we show that FIT is robust to distribution shifts over feature values and can generalize to new, unseen features, demonstrating its adaptability to dynamic user requirements and varying task contexts.
In sum, our primary contributions are as follows:
@@
    We introduce Focus Instruction Tuning (FIT), a method that allows users to flexibly and dynamically specify what features a model should or should not focus on when performing a task at inference time. FIT enables practitioners to incorporate domain-specific knowledge about core, spurious, or bias-relevant features in order to steer models according to the desired feature specification, whilst adding a level of explainability.
@@
    We show that FIT generalises with respect to (a) distribution shift over feature values and (b) focusing on new features not seen during training, further demonstrating Focus Instruction Tuning ’s adaptability.
Figure 1: Focus Instruction Tuning (FIT). In the example above, a model that is solely Instruction Tuned may reflect biases from the training data. For instance, in a question from BBQ (Parrish et al., 2022), when asked who posed a technical question at an engineering candidate’s interview involving both a man and a woman, the model might incorrectly answer ”the man” due to biases, despite evidence to the contrary. In contrast, a FIT model can ignore the gender feature and focus on the interview content, demonstrating steerability and adaptability at inference time.
@@ 2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset – i.e., input features that are correlated with outputs in the training distribution, but are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold Wang et al. (2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Xiao et al., 2021; Venkataramani et al., 2024; Arjovsky et al., 2019) or texture (Geirhos et al., 2018; Baker et al., 2018), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024; Borkan et al., 2019). For instance, the token SPIELBERG is spuriously correlated with positive sentiment in datasets like SST-2 (Socher et al., 2013b), meaning that models trained on SST-2 may learn to predict sentiment by leveraging these spurious features instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-causal features undermines the robustness of models in generalising to distribution shift.
A variety of approaches have been explored to detect and mitigate the effects of spurious feature learning, particularly under distribution shifts. Traditional approaches include prompt engineering (Sun et al., 2024), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), and directly incorporating causal inference strategies (Wang & Culotta, 2020; 2021; Udomcharoenchaikit et al., 2022). Substantial work in mechanistic interpretability has also aimed to discover models’ representation and use of task-causal or spurious features: for instance, causal probing (which trains probing classifiers to recognise and modify supervised feature representations encoded by foundation models; see Belinkov, 2022; Canby et al., 2024; Davies & Khakzar, 2024) has been used to study how models leverage causal versus spurious features features in the context of a given task (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating their use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, concept removal methods locate and manipulate supervised feature representations corresponding to bias features encoded by foundation models in order to remove these features (Ravfogel et al., 2020; 2022; 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
@@ 2.2 Controlling LLMs
Steering LLMs. While alignment techniques like Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022) have helped create more aligned and less harmful models, these models can still be prompted to generate harmful responses (Carlini et al., 2024). Furthermore, when these models are fine-tuned—currently the predominant paradigm for practical applications of LLMs—they have been shown to lose alignment with safety objectives and become more prone to producing harmful outputs (Qi et al., 2023).
This, combined with the practical infeasibility of fully retraining large foundation models, has driven the need for inference-time steering of LLMs to address safety misalignment. Recent work has proposed the use of steering vectors to modify attention weights (Bhattacharjee et al., 2024; Panickssery et al., ; Li et al., 2024) and the application of sparse autoencoders (SAEs) (Cunningham et al., 2023) to control the internal knowledge selection behaviour of LLMs (Zhao et al., 2024), enabling safer and more adaptive test-time responses. However, these methods typically assume white-box access to model representations, which can be complex and challenging to work with in practice. Furthermore, they do not train models to internalise test-time adaptability as a feature. Instead, they rely on external interventions for steering, lacking mechanisms for models to directly adapt their behaviour in response to user instructions provided in natural language.
Our work addresses these limitations by explicitly training models to exhibit controllable and robust test-time adaptability, enabling users to dynamically steer model behaviour through simple, natural language instructions.
@@ 3 Methodology
Let $\mathcal{F}$ denote the set of possible features (e.g., specific keywords, sentiment, verb tense, demographic information, etc.) that the model can be instructed to focus on or ignore when generating responses.
We consider a set of natural language instructions to focus or rule out specified features in $\mathcal{F}$ which we term the focus instruction set $\mathcal{I}_{\text{focus}}$. Explicitly, we define $\mathcal{I}_{\text{focus}}$ as
@@
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt in order to aid the model in learning the underlying task as well as the ability to refocus its attention on specified features during FIT. For specific examples of focus instrucitons that we consider, see Section A.4.
@@
FIT Training. Consider a classification task with finite label space $\mathcal{Y}$, where a single core feature $C\in\mathcal{F}$ is fully predictive of label $y\in\mathcal{Y}$ given input $x$ at both training time and under distribution shift (Koh et al., 2021). We also consider spurious features $S\in\mathcal{S}\subseteq\mathcal{F}$ from a subset of spurious features $\mathcal{S}$, where feature values $s\in\text{Val}(S)$ for some spurious feature $S\in\mathcal{S}$ correlate with a label $y_{s}\in\mathcal{Y}$, where this correlation may change under distribution shift (Ming et al., 2022). Finally, we define $\mathcal{F}$ as the set of features that may be included in focus instructions during training, consisting of the core feature and the set of spurious features =$\mathcal{F}=\{C\}\cup\mathcal{S}$.
For a sample $(x,y)\sim p_{\text{data}}(x,y)$, we specify the focus label $y_{\text{focus}}=y_{\text{focus}}(x,y,I_{\text{focus}})\in\mathcal{Y}$ that depends on the ground truth label $y$ and focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. Intuitively, we define focus label $y_{\text{focus}}$ as $y_{\text{focus}}=y$ when either no focus features are specified (i.e., using the empty focus instruction), when the focus is on the underlying core feature $C$, or when ignoring a spurious feature $S$; but when either the focus is on a spurious feature or the core feature is ignored, $y_{\text{focus}}$ is defined as the label spuriously correlated with a particular value of the spurious feature present in input $x$. This changing target $y_{\text{focus}}$ trains the model to learn to adjust its responses based on specified features. Formally, we define $y_{\text{focus}}$ as:
@@
(Defining $\rho_{\text{spurious}}$) . Let $S\in\mathcal{S}\subseteq\mathcal{F}$ denote a spurious feature. Suppose that a value of $S$, say $s$, is spuriously correlated with label $y_{s}$. Then we define
$\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s),$ (4)
where the $Y$ denotes a ground truth dataset label.
By varying $\rho_{\text{spurious}}(s)$, we can control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well, as core features.
To enable the model to effectively learn to utilise focus instructions and predict focus labels, thereby inducing steerable behaviour, we require that the spurious feature $S$ is not predictive of the ground truth label $Y$ during training, that $Y\perp\!\!\!\perp S$. Similarly, we require that the core feature $C$ is not predictive of the spurious label $Y_{S}$, that is $Y_{S}\perp\!\!\!\perp C$. To enforce these independence conditions within our synthetic datasets, described in Section 4, we set $\rho_{\text{spurious}}=1/N$ for tasks with $N$ class labels and ensure a balanced label distribution during training. We prove that these conditions are sufficient to enforce $Y\perp\!\!\!\perp S$ in our generated synthetic SS and SMNLI training datasets, which are further detailed in Section A.9 and Section A.11. A more comprehensive discussion of this approach, including the effects of varying $\rho_{\text{spurious}}$ during training, is provided in Section A.1.
@@
We further evaluate FIT under distribution shifts, where the specific values taken by spurious features do not overlap between the training and test sets, by introducing additional test sets:
[…]
We evaluate over these datasets specifically on our SMNLI dataset (c.f. Section 4.2).
@@ 4 Experiments
Although the primary focus of FIT is on adaptively steering LLMs at inference time, we include an additional debiasing experiment for comparison on the BBQ dataset in Section A.8 for completeness. While bias mitigation is a valuable and natural application of FIT, it is not its primary objective. Instead, this inclusion highlights FIT’s broader utility as a tool for model control and adaptability, demonstrating that it performs on par with dedicated bias mitigation techniques while also offering the unique advantage of test-time steerability.
@@
Baselines. We compare against in the main section of the paper: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline, $\text{SFT}(y_{\text{focus}})$, follows the same setup as the FIT method (trained on sampled inputs and focus labels), but without the inclusion of focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions in FIT. The few-shot baseline involves using 5 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. In Section A.6, we include two additional baselines: zero-shot and vanilla SFT for a more complete comparison with FIT. Further details of baselines and their results in comparison to FIT can be found in Section A.5.